40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jun 24 00:10:45 2012, yuta, Update, Green Locking, current beat electronics beatsetup20120623.png
    Reply  Wed Feb 6 14:51:25 2013, Jenne, Update, Electronics, 1 power supply replaced! 
Message ID: 6862     Entry time: Sun Jun 24 00:10:45 2012     Reply to this: 8008
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: current beat electronics 

I moved amplifiers for beat PD at PSL table to 1X2 rack. Current beat setup from PD to ADC is shown below. Setup for X beat and Y beat are almost the same except for minor difference like cable kind for the delay line.

Currently, DC power for amplifiers ZHL-1000LN+ is supplied by Aligent E3620A.
I tried to use power supply from the side of 1X1 rack, but fuse plug(Phoenix Contact ST-SI-UK-4) showed red LED, so I couldn't use it.
Measured amplification was +25 dB for 10-100 MHz.

Measured gain from RF input to monitor output of the beat box was ~ -1 db for 10-100 MHz.


ELOG V3.1.3-